Ferric oxide crystallises in a hexagonal close-packed array of oxide ions with two out of every three octahedral holes occupied by ferric ions. Derive the formula of the ferric oxide.

Grade:8

1 Answers

Sakshi
8 years ago
Sice ferric oxide crystallizes in hcp, thus, oxide ions occupy the lattice sites, i.e =6
ferric ions are present in octahedral voids, and the occupancy = 2/3
no of octahedral voids = 6
octahedral voids occupied = 2*6/3 = 4
thus, the formula of ferric oxide is
Fe 4 O6 or Fe2O3